OpenStudy (anonymous):

Suppose the diameter of a circle is 4. What is its area?

hero (hero):

The diameter is twice the radius. In other words, d = 2r. Therefore, r = 2. Use \[\text{Use} \space A = \pi r^2 \space \text {and solve.}\]

OpenStudy (anonymous):

A=3,14*4; A=12,56